OpenStudy (anonymous):

If f(x)=2+abs(x-3) for all x, what's the value of the derivative f'(x) at x=3? (Answer: Nonexistent)

OpenStudy (solomonzelman):

I've never a person with (x-3) abbs.

OpenStudy (anonymous):

abs means absolute value of something.

OpenStudy (shamil98):

\[f(x) = 2 + |x-3|\]

OpenStudy (anonymous):

That's right. Can you continue?

OpenStudy (shamil98):

Find the derivative, and then evaluate it at f'(3). I assume you need assistance finding the derivative , correct?

OpenStudy (anonymous):

Yes. Please find the derivative.

OpenStudy (shamil98):

Okay, let u = x-3 |u| = √u^2 \[u * \frac{ u' }{ |u| }\]

OpenStudy (shamil98):

sub back in (x-3)' = 1 u = x-3 |u| = |x-3| \[\huge \frac{ x-3 }{ |x-3| }\]

OpenStudy (shamil98):

I've never done derivatives involving absolute values, so I apologize in advance if this is incorrect, just my attempt.

OpenStudy (shamil98):

From what I've gathered, the basic rule for this is given through: The derivative of |x| is x/|x|. Applying the chain rule to this, the derivative of a function |u| is (u/|u|) * u'

OpenStudy (anonymous):

So (x-3)/(abs(x-3)) at x=3 is 0/0, which is 0, right?

OpenStudy (shamil98):

Ah, yes I forgot about the actual question, evaluating it at f'(3).

OpenStudy (shamil98):

0/0=undefined

OpenStudy (shamil98):

It is indeterminate, (i think that is the term).

OpenStudy (anonymous):

So undefined means nonexistent, right?

OpenStudy (shamil98):

Yes, it is non-existent, unless you were asked to solve for the indeterminacy .

OpenStudy (anonymous):

Thank you. You've helped a lot.